# Uncertainty Comes for Free: Human-in-the-Loop Policies with Diffusion Models

### v2

@@ Abstract
Keywords: Human-in-the-loop Policies, Policy fine-tuning
1 Introduction
@@
To achieve this, we utilize diffusion models [1] as our policy class. Diffusion policies offer two key advantages: (1) they have demonstrated robust performance in imitation learning tasks, and (2) their generative process involves an iterative denoising mechanism which we can leverage for insight into the agent’s decision-making process. Specifically, we use the denoising information to compute an uncertainty metric for the policy, which is then used during deployment to determine when human intervention is most beneficial (Fig. 1). To achieve this, we directly leverage the noise prediction model learned during the policy training process. Therefore, uncertainty estimation does not require training any additional models, imposes a minimal cost at run time, and can thus be considered a “free” byproduct of diffusion policy training. Finally, we show that data collected during the operator interventions can be incorporated back into training through a fine-tuning process that further improves policy performance. Our main contributions are as follows:
  • $\bullet$
@@
2 Related Work
Human-in-the-loop (HitL) approaches have been widely explored to enhance robot manipulation by incorporating various human feedback modalities, such as interventions [7, 8], preferences [9], rankings [10], scalar feedback [11], and gaze [12]. Recent works like HIL-SERL [13] and Sirius [14] show strong performance by leveraging human input during training. HitL methods are also common in autonomous driving, e.g., ZOOX’s re-routing system [15] allows operators to assist in challenging scenarios. However, most approaches rely on continuous human supervision and monitoring when assistance is actually needed during deployment. We propose an uncertainty-aware diffusion model that selectively requests human help in high-uncertainty states, reducing the need for constant oversight. Unlike HULA [16], which models uncertainty for online RL, our approach derives HitL policies from offline data, enabling more efficient and scalable deployment.
@@ 3 Method
Diffusion policies generate actions through an action-denoising process, leveraging denoising diffusion probabilistic models (DDPM). A DDPM models a continuous data distribution $p(a^{0})$ as reversing a forward noising process from $a^{0}$ to $a^{K}$, where $a^{K}$ is Gaussian noise sampled from $\mathcal{N}(0,\sigma^{2}\mathcal{I})$. The generative process $\pi(\mathrm{a_{t}}|\mathrm{o_{t}})$, where $\mathrm{o_{t}}$ and $\mathrm{a_{t}}$ are robot observations and actions at time step $t$, starts by sampling an action $a^{K}_{t}$ as random noise, and then iteratively denoises it using:
@@
3.1 Denoising-based uncertainty metric
@@
5:     if $\texttt{Uncertainty}(o_{t})\geq\texttt{threshold}$ then
@@
$D(\mathcal{V}^{s}_{t})=\frac{1}{n(n-1)}\sum_{i,j}1-S_{c}(g_{i},g_{j})\text{, where }S_{c}(g_{i},g_{j})=\frac{g_{i}\cdot g_{j}}{||g_{i}||\cdot||g_{j}||}$
@@
3.2 Uncertainty-based intervention and policy fine-tuning
@@
4 Experiments
Figure 2: Experiments in simulated environments. Left: we consider three scenarios during policy deployment. (a) Distribution shift; (b) Partial observability (c) Action multi-modality. Right: qualitative visualization of predicted uncertainty, with lighter colors indicating higher uncertainty.
@@ 4.1 Evaluation and Baselines
  • ThriftyDAgger [20], which uses a model-ensemble-based “novelty” (i.e., out-of-distribution) detection and a risk metric learned via Bellman updates on test-time data.
@@
We also try other metrics, namely Sentinel [21] and diffusion output variance, that can be potentially adapted into a HitL pipeline. We provide results and discussions about our attempts in Appendix A.
4.2 Simulated Environments
@@
We now summarize the simulated environments used to test our method; additional details about collected training data (e.g., observations) can be found in Appendix B. Distribution shift: Lift-sim. In this task, we ask the robot to grasp and lift objects in a table-top setting. To emulate distribution shift, demonstration data is collected using only a single object (red cube - see Fig.2), while for testing we roll out the pretrained policy to a set of unseen objects (round nuts, hammers, and hooks). Partial observability: Cup Stacking. Here, we ask the robot to grasp a green cup and place it inside a red cup. We use three views as our observation: front, side, and wrist. Successful execution requires the robot to infer object alignment based on its observations. Misalignment can lead to unintended collisions, leading to failures. To introduce variability, cup positions are randomized during data collection. Incorrect choice of action mode: Open drawer. Here, the robot is tasked with opening one of three drawers in the scene. The collected dataset includes trajectories for opening each drawer, with $1/3$ of the data corresponding to each drawer. However, the dataset does not specify which drawer is to be opened in a given trajectory, introducing under-specification.
@@
As a sanity check, we first evaluate the unassisted task performance of the diffusion policy on each task under the training data distribution. For the Lift-sim task, the fully autonomous policy achieves $100\%$ success rate on the training object but fails completely ($0\%$ success rate) on unseen objects. For Cup Stacking, the robot consistently picks up the first cup (100% grasp success rate) but fails to place it into the second cup due to alignment difficulties, resulting in a success rate of $0\%$ without human assistance. This task is also sensitive to observation selection – training with only side and front views causes the robot to fail when grasping the green cup. For Open Drawer, the fully autonomous policy learns to open a drawer with $100\%$ success if the task description does not specify which particular drawer should be opened. Interestingly, despite the under-specified training (i.e., no conditioning on which drawer to open), the policy captures the multi-modality of the training distribution. During 100 rollouts with random sampling (Fig. 10), the robot opens the middle and bottom drawers in 15% and 85% of trials, respectively, but never opens the top drawer.
4.3 Efficiency of Human Interactions
Lift- sim Cup- stacking Open- drawer
HULA-offline [16] 55.7($\pm$6.1) 54.0($\pm$16.3) 21.7($\pm$11.7)
ThriftyDAgger [20] 33.5($\pm$7.4) 21.2($\pm$15.6) 17.2($\pm$8.9)
Diff-DAgger [22] 30.2($\pm$1.3) 32.0($\pm$4.0) 16.0($\pm$4.4)
Our method 16.9($\pm$4.5) 5.4($\pm$1.0) 8.0(±($\pm$( ±1.9)
Avg. Full-traj. Length 76.6($\pm$5.9) 147.8($\pm$12.9) 114.8($\pm$5.7)
@@ 4.3 Efficiency of Human Interactions
Table 1: Average # of human assistance steps needed to achieve 100% success rates for simulated tasks.
Figure 3: Sensitivity to threshold selection.
@@
As shown in Table 1, for all simulated tasks, our method outperforms all baselines, and allows the policy to achieve perfect task success with the fewest intervention steps. Qualitatively (see accompanying video), we observe that, for Lift-sim, the robot only seeks human assistance when its gripper is close to the object, and lifting happens without intervention. For Cup Stacking, our method identifies states where the agent aligns the two cups as having high uncertainty, whereas picking up a cup (which benefits from unoccluded view) is marked as low uncertainty. Finally, for Open Drawer, the policy asks for assistance when it needs to decide which drawer to reach to, and, once the human operator steers it towards the intended target, the robot autonomously completes the rest of the task.
@@
4.4 Fine-tuning Performance
Figure 4: Average success rate of fine-tuning the Lift-sim task with different number of human intervention steps.
@@ 4.4 Fine-tuning Performance
4.5 Real robot experiments
Figure 5: Real robot experiments: we design our experiments to elicit the challenges described in Sec. 4 on a real robot. Examples of unassisted, HitL, fine-tuned policy rollouts can be found in the supplementary video .
Finally, we validate our method on real robot data collected via tele-operation. To support real-world deployment, we employ denoising diffusion implicit models (DDIM) [29] for high-frequency action generation. We evaluate our method on 4 real robot tasks (see Fig. 5). As in the simulated experiments, we show an example in each of the deployment problems. Additional implementation details of using DDIM with our method and detailed environment descriptions are provided in Appendix B.
@@ 4.5 Real robot experiments
Our method 7.2 6.8 6.5 8.4
Avg. Full-traj. Length 80.0 111.9 98.9 48.8
Table 2: Average # of human assisted steps needed to achieve 100% success rate during policy deployment.
@@
With HitL deployment, the robot can complete all four tasks. On average, our method only requests help from the human for approximately 8.3% of time steps for an 100% success rate (see Table 2). We note that, since we are using action chunking during real robot deployment, one human intervention allows the human to control the robot for four steps, the same as the diffusion policy. Details about the evaluations protocol (e.g., the number of rollouts for each evaluation) are included in Appendix B.
Qualitatively, our method identifies crucial states during policy execution. For example, in the Lift-real task, the robot asks for assistance when the gripper is close the the object. Using human-collected data with uncertainty, we can fine-tune the diffusion policy to improve 47% success rate on average (shown in Table 3), outperforming fine-tuning with full-trajectories of data. In the 3-Mode Pushing task, the robot autonomously reaches to the side of the object and then transfers control to the human operator, who poses the gripper in the correct location depending on the intended target.
Train Test $||\mathcal{D}_{ft}||$
Zero-shot 1 0.16 0
HitL fine-tuning (Ours) 1 0.63 80
Full-traj. fine-tuning 1 0.31 132
Table 3: Fine-tuning performance of the Lift-real task. Results are success rates derived by 20 policy rollouts per object.
5 Limitations
Our work investigates methods for seeking human assistance - we are primarily interested in automatically determining, at deployment time, when such assistance is needed, without overloading the operator with requests. However, we do not tackle other important aspects of a full HitL framework which we consider complementary to our work.
For example, we do not consider methods for conveying to the operator additional information regarding the sources of the uncertainty identified by the agent. In this work, there is no mechanism for identifying which deployment difficulty is affecting the task. We emphasize this as an under-explored challenge, given that uncertainty often stems from both training data and deployment-time conditions. Empirically, we observe that decomposing the denoising vector can sometimes offer insight into the source of uncertainty; for instance, high uncertainty in position versus rotation may point to different underlying issues. Another approach to reason about uncertainty source is combining our method with VLMs, which we see as our future work. Additionally, our method may fail when encountering extreme distribution shifts, such as a completely different background, where even human intervention cannot bring the robot back to a recoverable state for semi-autonomous execution.
Finally, we note that our method is limited by using diffusion-based policies. However, our method can be extended to flow-based policies [30, 31], which gains increasing attention in the field of robot learning due to its robust performance. In fact, flow-matching is equivalent to diffusion when using Gaussian as its source distribution [32]. We see extensions to flow-based methods as our future work.
6 Conclusions
In this work, we propose a novel method that enables robots to actively and efficiently request HitL assistance during deployment. Leveraging an uncertainty metric derived from the denoising process of diffusion policies, our approach identifies states where human intervention is most beneficial, thereby minimizing unnecessary monitoring and intervention. Experiments across diverse deployment scenarios demonstrate improved policy performance and adaptability, addressing a core challenge in HitL robotics: reducing human labor while enhancing autonomy and reliability. Furthermore, our method facilitates targeted data collection for policy refinement, showcasing its potential for interaction-driven learning.
For future work, our approach can be extended to Vision-Language-Action (VLA) models that employ diffusion-based action heads [33, 34]. We also aim to further automate human-robot interaction by exploring which types of information most effectively support communication. Specifically, we will investigate how to design interpretable feedback mechanisms that allow robots to express uncertainty and intent in a way that is intuitive for human operators—potentially leveraging Vision-Language Models, as demonstrated in [21]. These efforts aim to bridge the gap between fully autonomous systems and HitL deployment, fostering more efficient and scalable solutions for real-world robotic applications.
Acknowledgments
We appreciate hardware supports from Eric Chang, Sharfin Islam and Weizhe Ni. We also thank Katelyn Lee for providing experimental objects in the real robot experiments.
Appendix A Extra baselines
In this section, we also explore alternative methods that could potentially be adapted into a HitL pipeline. While promising in their original use cases, we found these methods unsuitable for online HitL settings. Therefore, we present our findings here rather than including them as baselines in our standardized comparisons. All evaluations are conducted on the Cup Stacking task.
Figure 6: Qualitative results using (a) Sentinel [21] and (b) output variance as metrics for HitL policy deployment.
Sentinel [21] is a method that utilizes action temporal consistency to analyze whether the agent is stopping to make progress. To compare with Sentinel, we train our diffusion policies with a longer action horizon of 8 steps, and use their Maximum Mean Discrepancy (MMD) with radial basis function (RBF) kernels to calculate in consistency. To derive best performance of Sentinel, we do a hyperparameter search on how many steps we should use to calculate MMD and select the best results based on whether it assigns the states that are visually occluded high MMD.
We observe that Sentinel performs poorly on our tasks for two main reasons. First, it often assigns high inconsistency to states involving reversed motions (e.g., approaching and lifting an object), even when the robot can complete these actions autonomously. This is due to its reliance solely on action inconsistency, which is naturally high in such transitions. Second, its out-of-distribution (OOD) detection suffers from higher latency because it depends on past actions for computing consistency. While this makes Sentinel effective for failure detection, it is ill-suited for HitL policies, where delayed intervention can result in unrecoverable failures (e.g., knocking over another cup in the Cup Stacking task).
Output Variance. We also experimented with directly using output variance as a metric for HitL deployment. However, this approach proved unstable and often produced false negatives. For instance, in the Cup Stacking task, the model assigns the highest variance to early trajectory steps (e.g., approaching the first cup), while showing low variance during critical failures (e.g., cup insertion). This occurs because early stages exhibit high single-mode variance, whereas during insertion, the model outputs fall into two distinct modes, reducing single-mode variance. These results underscore the importance of accounting for the multi-modality present in human demonstrations.
HULA-offline  implementation details. To make HULA-offline  compatible to other baselines, we augment all imitation learning datasets by labeling rewards if they are not available from the original environment: each step receives a reward of $1$ if it is the final step of the trajectory and 00 otherwise.
In this work, we extensively test the performance of our baselines to ensure fair comparisons. We find that the training of HULA-offline is sensitive to hyper-parameters. The underlying cause is the requirement of training a good Q-function in the base RL algorithm CQL. For example, Fig. 7 shows an example of training curves of one of our failed HULA training. Since the uncertainty estimation is dependent on the estimated Q function, the uncertainty estimation can be instable too. We also observe that HULA-offline  is more prone to over-fitting in both pretraining and fine-tuning than all other methods. It can complete the task in locations seen in the training set, but fails when the object locations are different, even if close to locations in the training set. This shows that our method can identify states that the policy needs more information.
Figure 7: Example of the training curve for an unstable training with HULA-offline.
Appendix B Training details, experiment setup and data statistics
B.1 Experiment set up
In this section, we detail the training and implementation setup for each experiment. The hyperparameters used during pretraining are provided in Table 4. We employ a transformer-based diffusion policy (DP-T) [1], which yields stable performance across the tasks tested. For each experiment, we use the best-performing policy from pretraining since policies with lower performance typically require more human assistance. Parameters for policy rollouts are also listed in Table 4.
@@ B.1 Experiment set up
Lift- sim Cup stacking Open drawer Lift- real Ramekin stacking 3-mode pushing Nut PnP
@@
Observations WFL SWFL FL Two side cameras views, one wrist view and low-dim obs
@@
##\## of eval rollouts 20 per object 20 15 20 per object 20 15 20
Table 4: Training details: S, W, F represents side, wrist, front camera views receptively. L represents low-dimensional observations that contains end-effector pose and gripper state.
B.2 Real world environment
In this work, human demonstrations are gathered using a tele-operation system (see Fig. 8) in which the robot is controlled by a trakSTAR electromagnetic 6-DoF pose tracker and a continuous gripper control unit. To successfully deploy the policies in the real world, we use DDIM for sampling actions to deploy the policy at 10Hz. Since DDIM is compatible with DDPM-based training, our method applies directly to DDIM models. Moreover, our vector field sampling can be parallelized and executed in a batched, feed-forward manner, resulting in minimal computational overhead during deployment. Here we describe the setting of each task presented in the main text.
Figure 8: Data collection pipeline.
Lift-real. This task extends the lift-sim setup to the real world, where the robot is trained to grasp a set of objects (Fig. 5) and tested on unseen objects to evaluate generalization. Although some objects from train and test set are visually similar, the robot needs to learn different strategies to grasp them. For example, the cup for testing has a higher radius than the one in the training set, making it difficult to cage with a gripper. The robot needs to learn to grip the rim of the cup to achieve stable grasps.
3-Mode Pushing. In this task, the robot is required to push a cup to three designated locations on the table, marked with blue tape. The dataset used for training includes equal proportions of pushing trajectories for each of the three locations, with each mode a third of the full dataset. Notably, during training, the specific target location (mode) for the push is not explicitly provided to the robot.
Figure 9: Examples of observations for Ramekin Stacking in different stages. The top row shows image frames from one of the side cameras and the bottom row shows the view of the wrist camera.
Ramekin Stacking. In this task, the robot is required to pick up one ramekin and place it on top of the other. The success criterion for this task is achieving stable placement, where the top ramekin is horizontally aligned with the bottom one. This requires precise alignment between the two ramekins. As shown in Fig. 9, a key challenge arises during stacking because the bottom ramekin becomes occluded in the wrist camera view, making it difficult for the robot to achieve proper alignment.
Nut Pick-and-Place. The final task evaluates our method in a pick-and-place scenario, where the robot must pick up a nut and place it onto a lug. Success is defined as a stable placement of the nut upon gripper release, ensuring proper alignment for threading. This task requires high placement precision, which is particularly challenging to achieve using visual observations alone.
B.3 Human-in-the-loop policy deployment
B.4 Steering multi-modal policy
A notable deployment challenge we highlight is the multi-modality of training data. We observe that diffusion policies typically recover two modes autonomously (see Fig. 10 and Fig. 11), despite the dataset containing three. However, when the robot is steered into key states via human assistance, it can recover and execute the third mode—suggesting that the policy has indeed learned all modes but requires appropriate initialization to access those not encountered during unassisted deployment. We note that in this scenario, the policy does not have to be uncertain – it knows how to execute all modes. However, it still shows inter-mode divergence since there are more than one options for it to execute.
Figure 10: Multi-modal demonstrations and learned modes from diffusion policy in the Open Drawer task.
@@ B.4 Steering multi-modal policy
Appendix C Hyperparameter exploration
In this section, we investigate how hyper-parameters affect the performance of our HitL agent. The choice of these hyper-parameters plays an important role of our uncertainty estimation, and hence can affect how the robot ask for human assistance.
Sampling Radius. The radius parameter $r$ defines the neighborhood size for collecting denoising vectors in uncertainty estimation. It is crucial to consider the scale of $r$ relative to the action distribution that a diffusion policy aims to recover. If $r$ is too large, the uncertainty metric becomes similar across all states, as it incorporates a broad range of them. Conversely, if $r$ is too small, the metric may be overly influenced by local uncertainty. In this work, we normalize each dimension of the action space to $[0,1]$, simplifying the selection of $r$.
In HitL fine-tuning for Lift-sim (see Table 6), we test radii ranging from 0.01 to 0.1. A radius of 0.05 achieves the best balance between intervention steps and success rate, requiring only 20 interventions while improving the success rate by 0.63. This efficiency results from accurate uncertainty detection when the gripper approaches unseen targets but fails to grasp them. Smaller radii and larger radii yield less precise estimations, leading to interventions that are either premature or delayed, reducing success rates despite more interventions.
Overall, a radius of 0.05 consistently achieves optimal performance by balancing local and global uncertainty estimations. Smaller radii miss key trajectory patterns, while larger radii incorporate irrelevant vectors, reducing the accuracy of uncertainty estimation in robotic manipulation tasks.

### v3

@@ Abstract
11footnotetext: indicates joint-first authorship.
I Introduction
@@
To achieve this, we utilize diffusion models [1] as our policy class. Diffusion policies offer two key advantages: (1) they have demonstrated robust performance in imitation learning tasks, and (2) their generative process involves an iterative denoising mechanism which we can leverage for insight into the agent’s decision-making process. Specifically, we use the denoising information to compute an uncertainty metric for the policy, which is then used during deployment to determine when human intervention is most beneficial (Fig. 1). To achieve this, we directly leverage the noise prediction model learned during the policy training process. Therefore, uncertainty estimation does not require training any additional models, imposes a minimal cost at run time, and can thus be considered a “free” byproduct of diffusion policy training.
A key feature of our metric is that it remains consistent across different tasks: we use the same fixed quantile-based threshold, estimated from a validation set, for all deployments, avoiding per-task calibration and underscoring the robustness and generality of our approach. Finally, we show that data collected during the operator interventions can be incorporated back into training through a fine-tuning process that further improves policy performance. Our main contributions are as follows:
@@
II Related Work
HitL approaches have been widely explored to enhance robot manipulation by incorporating various human feedback modalities, such as interventions [7, 8], preferences [9], rankings [10], scalar feedback [11], and gaze [12]. Recent works like HIL-SERL [13] and Sirius [14] show strong performance by leveraging human input during training. HitL methods are also common in autonomous driving, e.g., ZOOX’s re-routing system [15] allows operators to assist in challenging scenarios. However, most approaches rely on continuous human supervision and monitoring when assistance is actually needed during deployment. We propose an uncertainty-aware diffusion model that selectively requests human help in high-uncertainty states, reducing the need for constant oversight. Unlike HULA [16], which models uncertainty for online RL, our approach derives HitL policies from offline data, enabling more efficient and scalable deployment.
@@ III Method
[…]
@@
III-A Denoising-based uncertainty metric
@@
5:  if $\texttt{Uncertainty}(o_{t})\ge\texttt{threshold}$ then
@@
[…]
where,
[…]
@@
III-B Uncertainty-based intervention and policy fine-tuning
@@
IV Experiments
Figure 2: Experiments in simulated environments. We consider three scenarios during policy deployment. (a) Distribution shift; (b) Partial observability (c) Action multi-modality.
@@ IV-A Evaluation and Baselines
  • ThriftyDAgger [20], which uses a model-ensemble-based OOD detection and a risk metric learned via Bellman updates on test-time data.
@@
We also try other metrics, namely Sentinel [21] and diffusion output variance, that can be potentially adapted into a HitL pipeline. We summarize discussions about our attempts in Appendix -A.
Figure 3: Qualitative visualization of predicted uncertainty, with lighter colors indicating higher uncertainty
IV-B Simulated Environments
@@
We now summarize the simulated environments used to test our method; additional details about collected training data can be found in Table III. Distribution shift: Lift-sim. In this task, we ask the robot to grasp and lift objects in a table-top setting. To emulate distribution shift, demonstration data is collected using only a single object (red cube - see Fig.2), while for testing we roll out the pretrained policy to a set of unseen objects (round nuts, hammers, and hooks). Partial observability: Cup Stacking. Here, we ask the robot to grasp a green cup and place it inside a red cup. We use three views as our observation: front, side, and wrist. Successful execution requires the robot to infer object alignment based on its observations. Misalignment can lead to unintended collisions, leading to failures. To introduce variability, cup positions are randomized during data collection. Incorrect choice of action mode: Open drawer. Here, the robot is tasked with opening one of three drawers in the scene. The collected dataset includes trajectories for opening each drawer, with 131/3 of the data corresponding to each drawer. However, the dataset does not specify which drawer is to be opened in a given trajectory, introducing under-specification.
@@
As a sanity check, we first evaluate the unassisted task performance of the diffusion policy on each task under the training data distribution. For the Lift-sim task, the fully autonomous policy achieves $100\%$ success rate on the training object but fails completely ($0\%$ success rate) on unseen objects. For Cup Stacking, the robot consistently picks up the first cup (100% grasp success rate) but fails to place it into the second cup due to alignment difficulties, resulting in a success rate of $0\%$ without human assistance. This task is also sensitive to observation selection – training with only side and front views causes the robot to fail when grasping the green cup. For Open Drawer, the fully autonomous policy learns to open a drawer with $100\%$ success if the task description does not specify which particular drawer should be opened. Interestingly, despite the under-specified training (i.e., no conditioning on which drawer to open), the policy captures the multi-modality of the training distribution. During 100 rollouts with random sampling, the robot opens the middle and bottom drawers in 15% and 85% of trials, respectively, but never opens the top drawer.
IV-C Efficiency of Human Interactions
@@
TABLE I: Average # of human assistance steps needed to achieve 100% success rates for simulated tasks.
Lift-
sim
Cup-
stacking
Open-
drawer
HULA-offline [16] 55.7(6.1) 54.0(16.3) 21.7(11.7)
ThriftyDAgger [20] 33.5(7.4) 21.2(15.6) 17.2(8.9)
Diff-DAgger [22] 30.2(1.3) 32.0(4.0) 16.0(4.4)
Our method 16.9(4.5) 5.4(1.0) 8.0(($\pm$1.9)
Avg. Full-traj. Length 76.6(5.9) 147.8(12.9) 114.8(5.7)
@@
As shown in Table I, for all simulated tasks, our method outperforms all baselines, and allows the policy to achieve perfect task success with the fewest intervention steps. Qualitatively (see accompanying video), we observe that, for Lift-sim, the robot only seeks human assistance when its gripper is close to the object, and lifting happens without intervention. For Cup Stacking, our method identifies states where the agent aligns the two cups as having high uncertainty, whereas picking up a cup (which benefits from unoccluded view) is marked as low uncertainty. Finally, for Open Drawer, the policy asks for assistance when it needs to decide which drawer to reach to, and, once the human operator steers it towards the intended target, the robot autonomously completes the rest of the task.
@@
IV-D Fine-tuning Performance
Figure 4: Left: Average success rate of fine-tuning the Lift-sim task with different number of human intervention steps. Right: Sensitivity to threshold selection.
@@ IV-D Fine-tuning Performance
Figure 5: Real robot experiments: we design our experiments to elicit the challenges described in Sec. IV on a real robot.
IV-E Real robot experiments
Finally, we validate our method on real robot data collected via tele-operation. To support real-world deployment, we employ denoising diffusion implicit models (DDIM) [29] for high-frequency action generation. We evaluate our method on 4 real robot tasks (see Fig. 5). As in the simulated experiments, we show an example in each of the deployment problems. We employ a transformer-based diffusion policy (DP-T) [1], which yields stable performance across the tasks tested. For each experiment, we use the best-performing policy from pretraining since policies with lower performance typically require more human assistance.
TABLE II: Average # of human assisted steps needed to achieve 100% success rate during policy deployment.
@@ IV-E Real robot experiments
Ours 7.2 6.8 6.5 8.4
Full traj. 80.0 111.9 98.9 48.8
@@
With HitL deployment, the robot can complete all four tasks. On average, our method only requests help from the human for approximately 8.3% of time steps for an 100% success rate (see Table II). We note that, since we are using action chunking during real robot deployment, one human intervention allows the human to control the robot for four steps, the same as the diffusion policy. Details about the evaluations protocol are included in Table III.
TABLE III: Training details: S, W, F represents side, wrist, front camera views receptively. L represents low-dimensional observations that contains end-effector pose and gripper state.
@@
Lift-
sim
Cup
stacking
Open
drawer
Lift-
real
Ramekin
stacking
3-mode
pushing
Nut
PnP
@@
Observations WFL SWFL FL
Two side cameras views,
one wrist view and low-dim obs
@@
#\# of eval rollouts
20 3
20 15
20 4
20 15 20
Qualitatively, our method identifies crucial states during policy execution. For example, in the Lift-real task, the robot asks for assistance when the gripper is close the the object. Using human-collected data with uncertainty, we can fine-tune the diffusion policy to improve 47% success rate on average (shown in Table IV), outperforming fine-tuning with full-trajectories of data. In the 3-Mode Pushing task, the robot autonomously reaches to the side of the object and then transfers control to the human operator, who poses the gripper in the correct location depending on the intended target.
TABLE IV: Fine-tuning performance of the Lift-real task. Results are success rates derived by 20 policy rollouts per object.
Train Test $\mathcal{D}_{ft}||$
Zero-shot 1 0.16 0
HitL fine-tuning (Ours) 1 0.63 80
Full-traj. fine-tuning 1 0.31 132
@@
V Hyperparameter exploration
In this section, we investigate how hyper-parameters affect the performance of our HitL agent.
Sampling Radius. The radius parameter $r$ defines the neighborhood size for collecting denoising vectors in uncertainty estimation. It is crucial to consider the scale of $r$ relative to the action distribution that a diffusion policy aims to recover. If $r$ is too large, the uncertainty metric becomes similar across all states, as it incorporates a broad range of them. Conversely, if $r$ is too small, the metric may be overly influenced by local uncertainty. In this work, we normalize each dimension of the action space to (0,1[0,1], simplifying the selection of $r$.
In HitL fine-tuning for Lift-sim (see Table VI), we test radii ranging from 0.01 to 0.1. A radius of 0.05 achieves the best balance between intervention steps and success rate, requiring only 20 interventions while improving the success rate by 0.63. This efficiency results from accurate uncertainty detection when the gripper approaches unseen targets but fails to grasp them. Smaller radii and larger radii yield less precise estimations, leading to interventions that are either premature or delayed, reducing success rates despite more interventions. Overall, a radius of 0.05 consistently achieves optimal performance by balancing local and global uncertainty estimations. Smaller radii miss key trajectory patterns, while larger radii incorporate irrelevant vectors, reducing the accuracy of uncertainty estimation in robotic manipulation tasks.